OpenStudy (anonymous):

Write y = 8x as a logarithmic function. A. log8x = y B. logx8 = y C. log8y = x D. logy8 = x

Parth (parthkohli):

You most probably mean \(y = 8^x\). Remember that \(a = b^c\) is the same as \(\log _b a = c\).